A certain liquid X has a normal boiling point of 110.30°C and a boiling point elevation constant =Kb0.71·°C·kgmol−1. A solution is prepared by dissolving some barium hydroxide (BaOH2) in 350.g of X. This solution boils at 110.7°C. Calculate the mass of BaOH2 that was dissolved.
Answer #1

∆T = 110.7 - 110.30 = 0.4 °C

∆T = Kb * molality

= 0.4 = 0.71 * molality

Molality = 0.5634

Moles / mass of solvent ( in kg) = 0.5634

Moles / (350*10^-3) = 0.5634

Moles = 0.1972

Mass of Ba(OH)2 = mole * molar mass

= 0.1972* 171.34

= 33.79 grams

If you have any questions please comment

If you satisfied with the solution please rate it thanks